OpenStudy (anonymous):

In 1980, the nominal value of GDP was R15400 million, while in 2000, the nominal value of GDP was R17690 million. Using 1970 as the base year (i.e GDP deflator of 1970 = 100), the GDP deflator of 1980 was 136 and the GDP deflator of 2000 was 162. Using 1970s prices, we can say that real GDP between 1980 and 2000 has: A. Increased by approximately R2290 million B. Increased by approximately R400 million C. Decreased by approximately R 2300 million D. Cannot be answered with the given information E. Decreased by approximately R400 million

OpenStudy (anonymous):

E. Decreased by app. R400 M. First get the Real GDP of 1980 by dividing 1980's nominal GDP by 1980's price (that's, 15400M/136). Then, multiply the answer by the base year's deflator-- 100. You now have 1132.35 M as your 1980 real GDP. Now, compute for the 2000's real GDP (that's, 17690M/162.) Then divide that by 100, we now have 10919M. Get the difference.

OpenStudy (anonymous):

E. Decreased by approximately R400 million real gdp of 1980 = nominal gdp / gdp deflator.=15400/1.36=11323.49 real gdp of 2000 = nominal gdp / gdp deflator.=17690/1.62=10919.749 so gdp has decreased by 403.73M hence option E